20230605 ECD

Summary:

We lost much of the first half due to small but thick clouds passing over from the SE - it almost looked like they were forming near the site. When we opened, we reconfigured to PEPSI for a repeat of the SN observation. We observed it at CD2 + CD5,4,6, but noticed the TMS alignment laser was on and turned it off before repeating the observation at CDs 2+5. We reconfigured to LUCI for the PSF morphology test (JCG) and then to MODS to verify the field aberration data. We checked off these 3 DDT and Engineering observations, and will try to execute the LBC TMS/FPIA on-sky test and another set of MODS standards, plus ESM work, during the next 1.5 ECD nights. The full nightlog is here: 20230605 UT Nightlog

Preparations

  • Update the configuration files before taking the verification set of field aberration data (done)
  • Align the 3 SX TMS collimators if there is support for this (this was done by Tim and Bruce, thanks!)

Night Plan

Tasks:

  1. Field aberration verification dataset with LUCI
    1. from Doug's comment to the 20230527UT nightlog
Suggested Adjustment for SX: 
   Tip = -45", Delta IE = 45" 
   Tilt =    -10", Delta CA = 10" 
  1. PSF test (Juan Carlos)
  2. followup SN PEPSI
  3. MODS spectrophotometric standard (all grating modes if possible)
  4. MODS imaging standard (M92 or M71)
  5. Test of FPIA/TMS automation SW
  6. (LBC thermal cycling 4-6 hours, as a backup for cirrus/clouds, but requires <1.5" seeing)

Night Prep

  1. Update the "Offset Collimation" X,Y,RX,RY values in the SX collimation tables to account for the additional Tip=-45", Tilt=-10".
    1. ssh tcs@tcs1 and cd /lbt/data/config/tcs/PSF
    2. cp -p SXSMLUCIFERCollimation.dat SXSMLUCIFERCollimation.20230527a.dat - the file was modified last on 20230527.
      1. Current values are: X= -0.79200 Y= 5.101000 RX= -233.711 RY = 417.894
    3. Josh obtains the Pointing Offsets that correspond to M2 Tip/Tilt = -45"/-10". These are:
      • SX M2 Pointing offsets (X,Y,RX,RY) = (0.511, 2.298, -441.899, 98.2) BUT, these corrections are sending M2 to -RX limit (RX = -899").
      • Doug is online - Josh gave Z7=+8000nm to move off the negative limit.
    4. Updated the SXSMLUCIFERCollimation.dat "Offset Collimation" line, so now it reads:
      • Current values now are: X= -0.28100 Y= 7.399000 RX= -675.610 RY = 516.094 ... AND we are once again in the -RX limit.
    5. Now Josh is again sending M2 Z7 = 8000 nm and we're off the RX limit.
    6. Clear Coefficients and send M2 Z7 = -2000 nm. We're off the limits.
    7. Josh is giving delta IE = +45" and delta CA = +10".
  2. Check the Persson star field (BS9155 field list) for the distribution - Doug will look at the data coming in and the first 4-5 stars may be sufficient since they're well distributed.

Clear or some cirrus and good seeing:

  1. Field aberration verification dataset with LUCI
  2. PSF test (Juan Carlos)
  3. followup SN PEPSI
  4. MODS spectrophotometric standard (all grating modes if possible)
  5. MODS imaging standard (M92 or M71)
  6. Test of FPIA/TMS automation SW

Poorer seeing (but less than 1.4-1.5")

  1. follow-up SN PEPSI
  2. MODS spectrophotometric standard
  3. LBC eng during the 2nd half, but if not ready, then LUCI spectroscopy

Night Log

02:30 Josh has opened and is checking pointing and collimation. Over to Alex's google doc log: 20230605 UT Nightlog

-- OlgaKuhn - 02 Jun 2023
Topic revision: r4 - 05 Jun 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback